# Diego Marcondes 1122 $0$-0002-6087-4821    Mariana Feldman 11 0009-0002-1961-9172    Junior Barrera 11 $0$-0003-0439-0475

### v1

@@ 5 Application: Boundary recognition of digits with noise
As an example, we treat the problem of boundary recognition in digits with noise. We consider the USDMNN with two layers, and windows contained in the $3\times 3$ square trained with a training sample of $10$, and a validation sample of $10$, $56\times 56$ binary images. The training and validation samples are the same used in [11], the initial windows were considered as the five point cross centered at the origin, and the initial characteristic functions of a fixed window were those that minimized the training loss of its neighbor window visited before, except to the window on which they differ, where the characteristic function is initiated randomly. The algorithms were implemented in python and are available at https://github.com/MarianaFeldman/USDMM. The training was performed on a personal computer with processor 13th Gen Intel Core i7-1355U x 12 and 16 GB of ram.
@@
We consider the intersection of union (IoU) error that is more suitable for form or object detection tasks. The IoU error of $\psi_{\theta}$ in the training and validation sample is, respectively, {linenomath}
$L_{t}(\theta)=1-\frac{1}{10}\sum_{k=1}^{10}\frac{|Y_{k}\cap\psi_{\theta}(X_{k})|}{|Y_{k}\cup\psi_{\theta}(X_{k})|}$ $L_{v}(\theta)=1-\frac{1}{10}\sum_{k=1}^{10}\frac{|Y^{(v)}_{k}\cap\psi_{\theta}(X^{(v)}_{k})|}{|Y^{(v)}_{k}\cup\psi_{\theta}(X^{(v)}_{k})|}$
@@
that is, the mean proportion of pixels in $Y\cup\psi_{\theta}(X)$ that are not in $Y\cap\psi_{\theta}(X)$ among the sample points in the training and validation sample.
@@
The results are presented in Table 1 and Figure 2. We trained the two layer USDMNN with batch sizes $1$ and $10$, sampling $10$ neighbors in the SLGDA for the characteristic functions and considering all neighbors in the SLGDA for the windows. For the batch size $1$, we considered $50$ epochs to train the windows and $100$ epochs to train the characteristic functions, while for the batch size $10$ we considered $19$ and $50$ epochs, respectively. These batch sizes refer to the characteristic function SLGDA and in both cases we considered a batch size of $10$ for the window SLGDA. The best result was obtained with batch size $10$, although it took around ten times more to train. The SLGDA for the characteristic functions took an average of $17$ epochs to reach the minimum with batch size $1$, and an average of $35$ epochs with batch size $10$, what is expected since with batch size $1$ the characteristic function is actually updated $10$ times at each epoch.
@@
The time to train these USDMNN was significantly greater than that it took to train CDMNN in [11], where the best architecture had a total training time of less than one hour. Moreover, as opposed to what was observed with CDMNN, the best validation error was obtained with a batch size of $10$ and was actually lesser than the minimum validation error obtained in [11] before fine training ($0.066$). Furthermore, the minimum training loss attained via USDMNN ($0.037$) was lesser than the minimum training loss obtained in [11] ($0.049$). Therefore, the USDMNN have obtained similar empirical results as the CDMNN in [11] without any prior information, but at a greater computational cost.
@@
Table 1: Results of the two layer USDMNN trained with batch sizes $b=1$ and $b=10$ for the characteristic function SLGDA. We present the minimum training loss observed during training; the training and validation loss of the trained USDMNN; the algorithm total time and time until the minimum validation error; the number of epochs to the minimum validation error and the average number of epochs until the minimum training error with fixed windows; and the number of visited windows.
b Min. Train Learned Train Min Val. Total time (h) Time to min. (h) Epochs to min. (W) Mean Epochs to min (f) W visited
1 0.050 0.090 0.1571 5.84 4.81 42 17 589
10 0.037 0.044 0.060 62.39 37.95 12 35 252
@@
Figure 2: Result obtained after applying to the validation sample each layer of the USDMNN trained with a batch size of $10$.
6 Conclusion
In this paper, we proposed a hierarchical SLGDA to train USDMNN and illustrated it in a simple example. We obtained empirical results as good as that of [11], but much more computational resources were necessary, so high algorithmic complexity was the price to pay to obtain good results without considering prior information to design a specific CDMNN architecture.
@@ 6 Conclusion
The example in this paper is only an illustration of the method, and it is necessary to improve the algorithm implementation. We are currently working on an efficient implementation of the hierarchical SLGDA that, we believe, will significantly decrease the training time, so methods based on USDMNN may also be competitive with CDMNN from a computational complexity perspective. With a more efficient algorithm, it will be possible to fine train the USDMNN and better study the effect of settings such as the number of layers and the window size on the model performance. An efficient method to train DMNN without the need of strong prior information may help popularize methods based on them among practitioners which do not have strong knowledge in mathematical morphology to design specific CDMNN architectures.

### v2

@@ 5 Application: Boundary recognition of digits with noise
As an example, we treat the problem of boundary recognition of digits with noise. We consider the USDMNN with two layers, and windows contained in the $3\times 3$ square trained with a training sample of $10$, and a validation sample of $10$, $56\times 56$ binary images. The training and validation samples are the same used in [22] and the initial windows were considered as the five point cross centered at the origin. The initial characteristic functions for the first sequence of windows were randomly chosen, while the initial characteristic functions of a later sequence of windows were those that minimized the training error of its neighbor visited before, except for the window on which they differ, where the characteristic function is initiated randomly. The algorithms were implemented in python and are available at https://github.com/MarianaFeldman/USDMM. The training was performed on a personal computer with processor Intel Core i7-1355U x 12 and 16 GB of RAM.
@@
We consider the intersection of union (IoU) error that is more suitable for form or object detection tasks. The IoU error of $\psi_{\theta}$ in the training sample is {linenomath}
$L_{t}(\theta)=1-\frac{1}{10}\sum_{k=1}^{10}\frac{|Y_{k}\cap\psi_{\theta}(X_{k})|}{|Y_{k}\cup\psi_{\theta}(X_{k})|}$ (5)
@@
that is, the mean proportion of pixels in $Y\cup\psi_{\theta}(X)$ that are not in $Y\cap\psi_{\theta}(X)$ among the sample points in the training sample. The respective error on the validation sample is denoted by $L_{v}(\theta)$.
@@
The results are presented in Table 1 and Figure 2. We trained the two layer USDMNN with batch sizes $1$, $5$ and $10$, sampling $8$ neighbors in the SLDA for the characteristic functions and considering all neighbors in the SLDA for the windows. These batch sizes refer to the SLDA for the characteristic functions, and in all cases we considered a batch size of $10$ in the SLDA for the windows. We considered $50$ epochs to train the windows and $100$ epochs to train the characteristic functions. Each scenario was trained ten times, starting from distinct initial values of the characteristic functions. We present the minimum, mean and standard deviation of the results over the ten repetitions in Table 1.
The best results were obtained with batch size $10$, and it seems that, with the initial windows and characteristic functions we are considering, it is not possible to properly train with batch sizes $1$ and $5$. The training and validation error of the USDMNN trained with batch size $10$ did not vary a lot over the repetitions, so the algorithm is not sensible to the initial value for batch size $10$. Moreover, with batch size $10$, the SLDA for the characteristic functions took in average around $80$ epochs to reach the minimum in all repetitions, and the SLDA for the windows took an average of $26$ epochs to the minimum over the ten repetitions, although the variation was great and there was a repetition in which the minimum was achieved after only five epochs.
@@
The time it took to train the USDMNN was greater than that it took to train CDMNN in [22], but it can be decreased with a more sophisticated implementation of the algorithms. The minimum training and validation error of the trained operator attained with USDMNN ($0.032$ and $0.052$) was slightly greater than the respective minimum error obtained in [22] (around $0.029$ and $0.042$). Therefore, the USDMNN have obtained similar empirical results as the CDMNN in [22] without strong prior information.
@@
Table 1: Results of the two layer USDMNN trained with batch sizes $b=1$, $b=5$, and $b=10$ for the characteristic function SLDA. The results are in the form Minimum - Average (Standard Deviation) over the ten repetitions. We present the minimum training error observed during training; the training and validation error of the trained USDMNN; the algorithm total time and time until the minimum validation error; the number of epochs to the minimum validation error; and the average number of epochs until the minimum training error with fixed windows.
b Min. Train Learned Train Min Val. Total time (h) Time to min. (h) Epo. to min. (W) Mean Epo. to min (f)
1 0.442 - 0.57 (0.07) 0.597 - 0.677 (0.053) 0.183 - 0.238 (0.031) 113.4 - 135.6 (16.1) 11.6 - 45.7 (37.2) 4 - 16.8 (15.4) 47.8 - 49.0 (0.8)
5 0.421 - 0.49 (0.04) 0.492 - 0.611 (0.076) 0.127 - 0.155 (0.015) 78.5 - 91.9 (20.3) 13.648 - 32.478 (19.0) 4 - 18.7 (11.9) 51.4 - 52.9 (1.3)
10 0.031 - 0.036 (0.004) 0.032 - 0.039 (0.006) 0.052 - 0.056 (0.003) 173.6 - 182.6 (13.3) 14.262 - 98.324 (70.9) 5 - 26.2 (17.7) 80.9 - 82.6 (1.3)
@@
Figure 2: Result obtained after applying, to the validation sample, each layer of the USDMNN trained with a batch size of $10$ that had the least validation error.
6 Next steps and future research
In this paper, we proposed a hierarchical SLDA to train USDMNN and illustrated it in a simple example. We obtained empirical results almost as good as that of [22], but without the necessity to design a DMNN based on prior information. The next step of this work is to develop an efficient implementation of the algorithms and perform a more extensive experimental study to better understand features of them. In particular, it is necessary to better understand the sensitivity to the initial values and the role of the batch size on both the SLDA for the windows and for the characteristic functions. A more thorough comparison with the CDMNN proposed in [22] is also necessary.
@@ 6 Next steps and future research
We are currently working on an efficient implementation of the hierarchical SLDA that, we believe, will significantly decrease the training time, so methods based on USDMNN may also be competitive with CDMNN from a computational complexity perspective. With a more efficient algorithm, it will be possible to fine train USDMNN and better study the effect of settings such as the number of layers and the window size on the performance. An efficient method to train DMNN without the need of strong prior information may help popularize methods based on them among practitioners which do not have strong knowledge in mathematical morphology to design specific CDMNN architectures.
A promising line of research is to compare the USDMNN and CDMNN with other methods proposed in the literature. Since the DMNN are exact representations of morphological operators, an interesting study would be to compare them with other exact methods, rather than with approximations such as the classical MNN. In particular, it would be interesting to compare the DMNN with the binary morphological neural networks proposed in [1, 2]. We leave such a comparison as a topic for future research.